# Leveraging recurrence in neural network wavefunctions for large-scale simulations of Heisenberg antiferromagnets: the square lattice

### v2

@@ III.3 Do better variational energies guarantee better correlations?
Fig. 6 shows the relative error of the finite-size squared sublattice magnetization estimates plotted against the relative error of the variational energies for the SLAHM with periodic boundary conditions for $L$ up to 20, which is the largest system size for which we computed $M^{2}$. We compute the relative error of the squared sublattice magnetization with respect to the QMC estimates found in Ref. [10]. We compute the relative error of the variational energy with respect to QMC energies that will appear in a forthcoming manuscript [75]. This figure confirms that the most accurate energies also yield the most accurate estimates of the sublattice magnetization, which is consistent with the the results reported in the previous sections; however, below a certain threshold, there is not a strong trend between the two relative errors. Other VMC studies have examined the decoupling of the accuracy of correlation functions from the accuracy of the energy [34]. It has also been noted that achieving highly accurate correlations is more challenging than achieving high accuracy for the variational energy [88], but this does not fully explain the observed behavior.
@@ IV Conclusion
While the practical similarity between classical machine learning tasks and NQS is an active area of research, it is worth drawing connections between our results and relevant topics in the classical machine learning literature. For instance, a significant discovery in the modern age of deep learning is the existence of “neural scaling laws” which seem to generally describe how training loss behaves as a function of certain scaling variables such as the number of parameters in a neural network, the size of the dataset, or the amount of compute used during training [91]. Further investigation might reveal that the systematic improvement of our results could be quantified by a power law scaling of the training time used to obtain them. Notably, These power laws have also been identified for neural networks trained with transfer learning [92].
@@
It is also possible to look to machine learning literature for insights into the success of iterative retraining. It is understood that the recurrent nature of RNNs, which allows for weight sharing, is the property that permits iterative retraining; however, the more subtle question is why this technique is successful. Indeed, it is not always the case that neural networks trained to represent ground states are able to generalize in system size [93]. What is learned about the ground state of one finite-size system that is transferrable to the ground state of the next system size? The fact that our RNN wavefunctions generalize from one system size to the next suggests that they are “feature learning”, or learning non-trivial patterns in the data rather than memorizing the data. Features that would be transferrable between different system sizes could be e.g. correlations.
@@
Ultimately, this work serves to carefully benchmark the performance of iteratively retrained RNN wavefunctions for a well-understood problem. In Appendix G, we directly compare all of our estimated quantities with values from the literature, showing the success of our methods. Our results demonstrate that RNNs are a powerful tool capable of accurately representing ground states for finite-size systems and providing reliable estimates of ground-state observables in the thermodynamic limit. Having proved the effectiveness of these methods, the next step is to perform similar simulations for the more challenging triangular lattice spin-$\frac{1}{2}$ antiferromagnetic Heisenberg model (TLAHM).
@@ V Code Availability
Our implementation of the presented methods, all scripts needed to reproduce our results, and the data needed to reproduce the figures in this manuscript are openly available on GitHub https://github.com/mschuylermoss/HeisenbergRNN. Our code relies on TensorFlow [94], NumPy [95], and Matplotlib [96].
@@ Appendix A Marshall-Peierls Sign Rule
which rotates all spins on sublattice B by $\pi$ around the z-axis. Applying this unitary to our Hamiltonian defined in Eq. 1, we find that the off-diagonal components take on a leading factor of $-1$,
@@ Appendix B Details of RNN Wavefunctions
where $[\vec{a},\vec{b}]$ denotes the concatenation of two vectors. Note that for periodic boundary conditions, we have $\vec{h}_{i,L+1}\equiv\vec{h}_{i,1}$ and $\vec{h}_{L+1,j}\equiv\vec{h}_{1,j}$ as well as $\vec{\sigma}_{i,L+1}\equiv\vec{\sigma}_{i,1}$ and $\vec{\sigma}_{L+1,j}\equiv\vec{\sigma}_{1,j}$.
For our simulations of systems with open boundary conditions, which can be found in Appendix F, we use a Tensorized version of the GRU cell, which improves the expressiveness of the ansatz [99] at the cost of a larger number of parameters. This cell only accepts the two hidden vectors and spin states that come from the nearest-neighbor cells that appear previously in the autoregressive sequence. The candidate hidden vector and update gate are computed as follows:
@@ Appendix D Runtime Details
Figure 8: The amount of time in seconds it takes to complete the training up to a given system size with periodic boundary conditions using four A100 GPUs. The solid lines show an estimate of the training time if the number of training steps used for $L=6$ is used for all $L$.
@@
In practice, one expects the optimization of a neural network quantum state to become more challenging for larger system sizes if one is optimizing from scratch, due to the larger Hilbert space. If this is the case, then for large simulations one would need to train longer in order to converge to the ground state. As a more generous comparison, we show how the run time would scale if the number of training steps is held constant at $N_{\mathrm{steps}}=s\times C$ for all lattice sizes (still keeping the network size fixed). Even if this would be sufficient, our training schedule reduces the run time by an order of magnitude or more compared to this baseline. We do not study how the choice of $d_{h}$ should scale with $N$, so we are unable to comment on how the run time would scale if RNN wavefunctions of different size were used for physical systems of different size.
@@ Appendix E V-Score Scaling
Appendix F Results for the SLAHM with open boundary conditions
@@
Fig. 11(a) shows that our variational energies are in very close agreement with reference values obtained from gradient-optimized PEPS simulations [100]. In some cases, our variational energies are lower than the reported PEPS values, which is consistent with previous works [38]. Unlike what we observed with the simulations of the SLAHM with periodic boundary conditions, none of the variational energies strongly diverge from the reference values. Fig. 11(b) displays the variances of our variational energies. Although the variational energies seem to be in good agreement across all simulations, the variances still improve systematically based on the training schedule. Training with a larger scale $s$ or a slower rate $r$, and thus lengthening the total training time, improves the variance of the variational energy. Interestingly, these variances also hint at some instability in the training for larger system sizes when the scale $s$ is smaller. This suggests that, at least in the case of open boundary conditions, the RNN wavefunctions for a given system size $L$ must be sufficiently trained before moving on the the next system size. Only the three simulations with the longest training times maintain low variances across all system sizes. In general, these variances are much lower than the variances obtained from our simulations of the SLAHM with periodic boundary conditions.
@@
Figure 13: Estimates of the squared sublattice magnetization scaled as a function of $1/L$. These estimates of $M^{2}$ and $M_{z}^{2}$ were obtained from our simulations with the longest amount of training time, i.e. $s=4.0$ and $r=0.25$. Closed markers represent the values of $M^{2}$ and $M_{C}^{2}$ estimated using the correlations defined by Eq. 8, while open markers represent the values of $M^{2}$ and $M^{2}_{C}$ that are estimated using the correlations defined by Eq. 9. The thermodynamic limit (TL) value of the squared sublattice magnetization from QMC are shown for reference.
We also examine estimates of the sublattice magnetization obtained from our simulations of open systems with $L$ up to 20. In particular, we show estimates of the sublattice magnetization obtained from the structure factor, as defined in Eq. 11. We show both estimates of this value: one obtained using the full correlations defined in Eq. 8 and one obtained using only the z-basis correlations defined in Eq. 9. Fig. 13 shows our estimates of these values obtained from our longest and most accurate simulation, i.e. with the largest scale $s=4.0$ and the slowest rate $r=0.25$. We also perform a finite-size scaling of these values using a second-order polynomial in $1/L$. Liu et al. [100] performed a similar finite-size scaling, but only reference values for their extrapolated value of $M^{2}$ were provided. The extrapolated value obtained from our finite-size scaling is reduced from the value obtained from the PEPS simulation, and reduced from the value obtained from QMC [10]. This is likely due to finite-size effects which are stronger for systems with open boundary conditions. Indeed, the extrapolated value of $M_{\infty}^{\text{PEPS}}$ was obtained from values of $M^{2}$ that were estimated by considering only the smaller, central portion of the lattice precisely to reduce edge effects.
@@
Appendix G Comparison of results to benchmarks and other methods
@@
periodic RBM [21, 42] $-0.668(5)$
$6\times 6$ CNN [33] $-0.67882(1)$
RBM+Lanczos [40] $-0.678868(2)$
@@
periodic CNN [33] $-0.67135$
$10\times 10$ RBM+Lanczos [40] $-0.671519(4)$
@@
open PixelCNN [36] $-0.628627(1)$
$10\times 10$ RNN [38] $-0.628638(4)$
@@
Spin-Wave Theory [5, 6, 7] $-0.6705$ 0.303
Series expansions [101] $-0.6696(3)$ 0.313
Exact Diagonalization [102] $-0.672(1)$ 0.25
QMC - World-line [4] $-0.670(1)$ 0.31
QMC - Green’s function [103] $-0.6692(2)$ 0.31(2)
QMC - Green’s function [104] $-0.6692(2)$ 0.34(1)
QMC - Stochastic Series Expansion [9] $-0.669437(5)$ 0.3070(3)
@@
VMC - RVB ansatz [15] $-0.6682(4)$
VMC - Gutzwiller projector ansatz [13] $-0.642$ $\sim$0.43
VMC - Jastrow ansatz [14] $-0.6638$ $\sim$0.4
VMC - Marshall-Jastrow ansatz [16] $-0.6637(2)$ 0.349(2)
@@
gradient optimized iPEPS [106] $-0.669432$ 0.3152
gradient optimized PEPS [100] $-0.66948(42)$ 0.305

### v3

@@ III.3 Do better variational energies guarantee better correlations?
Fig. 6 shows the relative error of the finite-size squared sublattice magnetization estimates plotted against the relative error of the variational energies for the SLAHM with periodic boundary conditions for $L$ up to 20, which is the largest system size for which we computed $M^{2}$. We compute the relative error of the squared sublattice magnetization with respect to the QMC estimates found in Ref. [10]. We compute the relative error of the variational energy with respect to QMC energies that will appear in a forthcoming manuscript [76]. This figure confirms that the most accurate energies also yield the most accurate estimates of the sublattice magnetization, which is consistent with the results reported in the previous sections; however, below a certain threshold, there is not a strong trend between the two relative errors. Other VMC studies have examined the decoupling of the accuracy of correlation functions from the accuracy of the energy [34]. It has also been noted that achieving highly accurate correlations is more challenging than achieving high accuracy for the variational energy [89], but this does not fully explain the observed behavior.
@@ IV Conclusion
While the practical similarity between classical machine learning tasks and NQS is an active area of research, it is worth drawing connections between our results and relevant topics in the classical machine learning literature. For instance, a significant discovery in the modern age of deep learning is the existence of “neural scaling laws” which seem to generally describe how training loss behaves as a function of certain scaling variables such as the number of parameters in a neural network, the size of the dataset, or the amount of compute used during training [92, 93, 94]. Further investigation might reveal that the systematic improvement of our results could be quantified by a power law scaling of the training time used to obtain them. Recently, neural scaling laws have been observed in the context of NQS applied to quantum chemistry problems [95]. Ref. [95] demonstrates a power law decay in ground-state energy accuracy, using the number of parameters in their ansätze as the main scaling variable.
In the present work, we considered only one network size, $d_{h}=256$, which we chose based on previous iterative retraining studies [35, 38]. Since our RNN wavefunctions show systematic improvement with increased training time, we believe that the quality of our simulations is not limited by the expressiveness of the ansätze. Still, the size of the hidden vectors $d_{h}$ is an important hyperparameter to explore further. For one, the scaling of $d_{h}$ could be the key to revealing the aforementioned neural scaling laws. Second, our iterative retraining experiments would provide an interesting way to study how $d_{h}$ should scale with the size of the system $N=L^{2}$. If an RNN wavefunction is not expressive enough to accurately represent the ground state for a given system size $L$, then increasing the training time should not improve the accuracy. Furthermore, increasing $L$ beyond the system size where accuracy saturates due to limitations in expressiveness should lead to an observable degradation in simulation quality. Given the connections between matrix product states and RNN wavefunctions [41, 96], it is possible that there exists a physically meaningful way to scale $d_{h}$ with system size $N$. We leave such experiments for future work.
@@
It is also possible to look to machine learning literature for insights into the success of iterative retraining. It is understood that the recurrent nature of RNNs, which allows for weight sharing, is the property that permits iterative retraining; however, the more subtle question is why this technique is successful. Indeed, it is not always the case that neural networks trained to represent ground states are able to generalize in system size [97]. What is learned about the ground state of one finite-size system that is transferrable to the ground state of the next system size? The fact that our RNN wavefunctions generalize from one system size to the next suggests that they are “feature learning”, or learning non-trivial patterns in the data rather than memorizing the data. Features that would be transferrable between different system sizes could be e.g. correlations. The generalization properties of NQS are the focus of ongoing research [98, 99].
@@
Ultimately, this work serves to carefully benchmark the performance of iteratively retrained RNN wavefunctions for a well-understood problem. In Appendix H, we directly compare all of our estimated quantities with values from the literature, showing the success of our methods. Our results demonstrate that RNNs are a powerful tool capable of accurately representing ground states for finite-size systems and providing reliable estimates of ground-state observables in the thermodynamic limit. Having proved the effectiveness of these methods, the next step is to perform similar simulations for the more challenging triangular lattice spin-$\frac{1}{2}$ antiferromagnetic Heisenberg model (TLAHM) [100].
@@ V Code Availability
Our implementation of the presented methods, all scripts needed to reproduce our results, and the data needed to reproduce the figures in this manuscript are openly available on GitHub https://github.com/mschuylermoss/HeisenbergRNN [101]. Our code relies on TensorFlow [102], NumPy [103], and Matplotlib [104].
@@ Appendix A Marshall-Peierls Sign Rule
which rotates all spins on sublattice B by $\pi$ around the z-axis. Applying this unitary to our Hamiltonian defined in Eq. 1, we find that the off-diagonal components take on a leading factor of 1-1,
@@ Appendix B Details of RNN Wavefunctions
where $\vec{a},\vec{b}]$ denotes the concatenation of two vectors. Note that for periodic boundary conditions, we have $\vec{h}_{i,L+1}\equiv\vec{h}_{i,1}$ and $\vec{h}_{L+1,j}\equiv\vec{h}_{1,j}$ as well as $\vec{\sigma}_{i,L+1}\equiv\vec{\sigma}_{i,1}$ and $\vec{\sigma}_{L+1,j}\equiv\vec{\sigma}_{1,j}$.
For our simulations of systems with open boundary conditions, which can be found in Appendix G, we use a Tensorized version of the GRU cell, which improves the expressiveness of the ansatz [107] at the cost of a larger number of parameters. This cell only accepts the two hidden vectors and spin states that come from the nearest-neighbor cells that appear previously in the autoregressive sequence. The candidate hidden vector and update gate are computed as follows:
@@ Appendix D Runtime Details
Figure 8: The amount of time in seconds it takes to complete the training up to a given system size with periodic boundary conditions using four A100 GPUs.
@@
In practice, one expects the optimization of a neural network quantum state to become more challenging for larger physical system sizes if one is optimizing from scratch, due to the larger Hilbert space. If this is the case, then for simulations of large systems one would likely need to train longer (for more training steps) in order to converge to a good approximation of the ground state. Given that the time per training step grows rapidly with system size, as seen in Fig. 7, increasing the number of steps with $L$ would lead to a rapid increase in total run time. Iterative retraining and the ability to decrease the number of training steps as the system size $L$, and consequently the time per training step, increases, allow us to dramatically reduce the amount of computation time needed to obtain simulation results for many large system sizes.
Importantly, we do not study how the choice of $d_{h}$ should scale with $N$, so we are unable to comment on how the run time would scale if RNN wavefunctions of different size were used for physical systems of different size.
@@ Appendix E V-Score Scaling
Appendix F Zero-variance extrapolation of the variational energy
Using a sequence of systematically improving variational states, it is common to extrapolate the corresponding variational energies to their zero-variance limit [32, 81, 82, 83, 84, 85]. This type of extrapolation is motivated by the zero-variance principle [54]. Fig. 11 shows an example of this type of extrapolation for our best variational energies for $L=32$ and periodic boundary conditions. As mentioned, we perform this fit only through our best variational energies, which come from simulations with $s\geq 1.0$.
Figure 11: The zero-variance extrapolation of our best variational energies from Fig. 3 for $L=32$. Each variational energy comes from an RNN wavefunction trained according to the schedule given by Eq. 6 and a unique choice of the scale $s$ and rate $r$. The dotted grey line shows the linear fit through the data.
Appendix G Results for the SLAHM with open boundary conditions
@@
Fig. 12(a) shows that our variational energies are in very close agreement with reference values obtained from gradient-optimized PEPS simulations [108]. In some cases, our variational energies are lower than the reported PEPS values, which is consistent with previous works [38]. Unlike what we observed with the simulations of the SLAHM with periodic boundary conditions, none of the variational energies strongly diverge from the reference values. Fig. 12(b) displays the variances of our variational energies. Although the variational energies seem to be in good agreement across all simulations, the variances still improve systematically based on the training schedule. Training with a larger scale $s$ or a slower rate $r$, and thus lengthening the total training time, improves the variance of the variational energy. Interestingly, these variances also hint at some instability in the training for larger system sizes when the scale $s$ is smaller. This suggests that, at least in the case of open boundary conditions, the RNN wavefunctions for a given system size $L$ must be sufficiently trained before moving on the next system size. Only the three simulations with the longest training times maintain low variances across all system sizes. In general, these variances are much lower than the variances obtained from our simulations of the SLAHM with periodic boundary conditions.
@@
Figure 14: Estimates of the squared sublattice magnetization scaled as a function of $1/L$. These estimates of $M^{2}$ and $M_{z}^{2}$ were obtained from our simulations with the longest amount of training time, i.e. $s=4.0$ and $r=0.25$. Closed markers represent the values of $M^{2}$ and $M_{C}^{2}$ estimated using the correlations defined by Eq. 8, while open markers represent the values of $M^{2}$ and $M^{2}_{C}$ that are estimated using the correlations defined by Eq. 9. The thermodynamic limit (TL) value of the squared sublattice magnetization obtained from PEPS is shown for reference.
We also examine estimates of the sublattice magnetization obtained from our simulations of open systems with $L$ up to 20. In particular, we show estimates of the sublattice magnetization obtained from the structure factor, as defined in Eq. 11. We show both estimates of this value: one obtained using the full correlations defined in Eq. 8 and one obtained using only the $z$ basis correlations defined in Eq. 9. Fig. 14 shows our estimates of these values obtained from our longest and most accurate simulation, i.e. with the largest scale $s=4.0$ and the slowest rate $r=0.25$. We also perform a finite-size scaling of these values using a second-order polynomial in $1/L$. Liu et al. [108] performed a similar finite-size scaling, but only reference values for their extrapolated value of $M^{2}$ were provided. The extrapolated value obtained from our finite-size scaling is reduced from the value obtained from the PEPS simulation, and reduced from the value obtained from QMC [10]. This is likely due to finite-size effects which are stronger for systems with open boundary conditions. Indeed, the extrapolated value of $M_{\infty}^{\text{PEPS}}$ was obtained from values of $M^{2}$ that were estimated by considering only the smaller, central portion of the lattice precisely to reduce edge effects.
@@
Appendix H Comparison of results to benchmarks and other methods
@@
periodic RBM [21, 42] 0.668(5)-0.668(5)
$6\times 6$ CNN [33] 0.67882(1)-0.67882(1)
RBM+Lanczos [40] 0.678868(2)-0.678868(2)
@@
periodic CNN [33] 0.67135-0.67135
$10\times 10$ RBM+Lanczos [40] 0.671519(4)-0.671519(4)
@@
open PixelCNN [36] 0.628627(1)-0.628627(1)
$10\times 10$ RNN [38] 0.628638(4)-0.628638(4)
@@
Spin-Wave Theory [5, 6, 7] 0.6705-0.6705 0.303
Series expansions [109] 0.6696(3)-0.6696(3) 0.313
Exact Diagonalization [110] 0.672(1)-0.672(1) 0.25
QMC - World-line [4] 0.670(1)-0.670(1) 0.31
QMC - Green’s function [111] 0.6692(2)-0.6692(2) 0.31(2)
QMC - Green’s function [112] 0.6692(2)-0.6692(2) 0.34(1)
QMC - Stochastic Series Expansion [9] 0.669437(5)-0.669437(5) 0.3070(3)
@@
VMC - RVB ansatz [15] 0.6682(4)-0.6682(4)
VMC - Gutzwiller projector ansatz [13] 0.642-0.642 $\sim$0.43
VMC - Jastrow ansatz [14] 0.6638-0.6638 $\sim$0.4
VMC - Marshall-Jastrow ansatz [16] 0.6637(2)-0.6637(2) 0.349(2)
@@
gradient optimized iPEPS [114] 0.669432-0.669432 0.3152
gradient optimized PEPS [108] 0.66948(42)-0.66948(42) 0.305